There is a line through the origin that divides the region bounded by the parabola 

y = 6x - 7x2 and the x-axis into two regions with equal area. What is the slope of that line?


1
Expert's answer
2022-04-29T10:18:46-0400
"6x-7x^2=0"

"x_1=0, x_2=\dfrac{6}{7}"

"Area_{total}=\displaystyle\int_{0}^{6/7}(6x-7x^2)dx"

"=[3x^2-\dfrac{7}{3}x^3]\begin{matrix}
 6/7 \\
 0
\end{matrix}=\dfrac{36}{49}({units}^2)"

"A_1=A_2=\dfrac{18}{49}{units}^2"

Let "m=" the slope of a line: "y=mx"


"6x-7x^2=mx"

"x_1=0, x_2=\dfrac{6-m}{7}"

"Area_1=\displaystyle\int_{0}^{(6-m)/7}(6x-7x^2-mx)dx"

"=[3x^2-\dfrac{7}{3}x^3-\dfrac{m}{2}x^2]\begin{matrix}
 (6-m)/7 \\
 0
\end{matrix}"

"=\dfrac{3(6-m)^2}{49}-\dfrac{(6-m)^3}{3(49)}-\dfrac{m(6-m)^2}{2(49)}=\dfrac{18}{49}"

"(6-m)^2(18-12+2m-3m)=108"

"6-m=\sqrt[3]{108}"

"m=6-3\sqrt[3]{4}"